the sum of two numbers is 26. the difference of the numbers is 8. what are the numbers?

Found 3 solutions by stanbon, lwsshak3, solver91311:
Answer by stanbon(75887)

the sum of two numbers is 26. the difference of the numbers is 8. what are the numbers?
-----
Equations:
a + b = 26
a - b = 8
----
Add and solve for "a":
2a = 34
a = 17
-----
Solve for "b":
a + b = 26
17 + b = 26
b = 9
===========
Cheers,
Stan H.
============

Answer by lwsshak3(11628)

the sum of two numbers is 26. the difference of the numbers is 8. what are the numbers?
**
let x=one of the numbers
26-x=other number
..
x-(26-x)=8
x-26+x=8
2x=34
x=17
26-x=9
the numbers are 17 and 9

Answer by solver91311(24713)







Solve the 2X2 system for and

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism